OpenStudy (anonymous):

Find the 27th term in the sequence. 59, 56, 53, 50, ...

OpenStudy (cwrw238):

this is an arithmetic series with common difference -3

OpenStudy (cwrw238):

nth term = a1 + (n - 1)d

OpenStudy (anonymous):

a1=56,d=56-59=-3 n=27 find n27 by the above formula.